OpenStudy (anonymous):

Write An Equation Of The Line Containing The Points (3,2) And (0,-3). Use The For y=mx+b. A. y=-3/5x-3 B. y=3/5x+3 C. y=5/3x-3 D. y=5/3x-3

OpenStudy (anonymous):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] do some little algebra and eventually you will have slope - intercept form \[y=mx+b\]

OpenStudy (anonymous):

I Dont Understand Any Of That...